OpenStudy (anonymous):

show that the sequence {sn} is convergent sn=1.3.5.7...(2n-1)/ 2.4.6.8...2n is convergent pls help

OpenStudy (loser66):

Let generalize it \(S_n = \dfrac{2n-1}{2n}=\dfrac{2n(1-1/2n}{2n}= 1-1/2n\) Hence \(lim S_n= 1\) --> \(S_n \) converges